OpenStudy (nerdgamer_):

Helena made trail mix and wants to list the ingredients in order of increasing weight. She used one and five eighths lb of mixed nuts, five eighths lb of chocolate covered raisins, and one and two eighthslb of dried mixed fruit. If she decides to include three eighths lb of sunflower seeds, which list shows the order of ingredients from least weight to greatest weight? raisins, seeds, fruit, nuts seeds, raisins, fruit, nuts raisins, seeds, nuts, fruit seeds, raisins, nuts, fruit

OpenStudy (nerdgamer_):

medal to whoever helps me and a fan

OpenStudy (anonymous):

Seeds:3/8 Raisins:5/8 Fruit:1 2/8 = 10/8 Nuts:1 5/8 = 13/8

OpenStudy (anonymous):

seeds, raisins, fruit, then nuts

OpenStudy (nerdgamer_):

ok

OpenStudy (anonymous):

All you have to do is convert all of them into improper fractions

OpenStudy (nerdgamer_):

ohh ok

OpenStudy (nerdgamer_):

thankyou i give you a medal and fan ok

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

your welcome